# PACMANN: Point Adaptive Collocation Method for Artificial Neural Networks

### v1

@@ Appendix A Relative error behavior of all optimization algorithms considered
This appendix contains figures which describe the behavior of PACMANN in combination with all optimization methods listed in Section 2. Figure 8(a) illustrates the prediction accuracy of various optimization algorithms across a range of 500 to 20,000 collocation points in the Burgers’ equation example. Overall, the majority of algorithms demonstrate equivalent performance in terms of accuracy. Notably, the golden section search algorithm achieves the lowest error at large numbers of collocation points (20,000 points). Furthermore, we evaluate the behavior of the optimization algorithms for a varying resampling period, from 25 to 1000 iterations, as depicted in Figure 8(b). Most optimization algorithms exhibit a behavior similar to Adam, as highlighted in Section 3. Notably, PACMANN loses prediction accuracy slower for increasing resampling period compared to the other sampling methods.
Similarly, we vary the same hyperparameters for the Allen-Cahn equation example. As shown in Figure 9(a), the other optimization algorithms display behavior comparable to Adam, with prediction accuracy remaining largely unaffected by an increasing number of collocation points. A similar trend is observed for the resampling period in Figure 9(b), where the accuracy shows minimal variation as the number of iterations increases.

### v2

@@ Appendix A Relative error behavior of all optimization algorithms considered
This appendix contains figures which describe the behavior of PACMANN in combination with all optimization methods listed in Section 2. Figure 10(a) illustrates the prediction accuracy of various optimization algorithms across a range of 500 to 20 000 collocation points for the Burgers’ equation example in Section 3.1. Overall, the majority of algorithms demonstrate equivalent performance in terms of accuracy. Notably, the golden section search algorithm achieves the lowest error at large numbers of collocation points (20 000 points). Furthermore, we evaluate the behavior of the optimization algorithms for a varying resampling period, from 25 to 1000 iterations, as depicted in Figure 10(b). Most optimization algorithms exhibit a behavior similar to Adam, as highlighted in Section 3. Notably, PACMANN loses prediction accuracy slower for increasing resampling period compared to the other sampling methods.
Similarly, we vary the same hyperparameters for the Allen-Cahn equation example in Section 3.2. As shown in Figure 11(a), the other optimization algorithms display behavior comparable to Adam, with prediction accuracy remaining largely unaffected by an increasing number of collocation points. A similar trend is observed for the resampling period in Figure 11(b), where the accuracy shows minimal variation as the number of iterations increases.